### v1

# Infinitely many pairs of non-isomorphic elliptic curves sharing the same BSD invariants

Abstract.
We prove that there exist infinitely many pairs of non-isomorphic elliptic curves over $\mathbb{Q}$ sharing the same BSD invariants—including their $L$-functions, Mordell–Weil groups, Tate–Shafarevich groups, Tamagawa numbers, regulators, and real periods—and Kodaira symbols and minimal discriminants, while having distinct $j$-invariants.
1. Introduction
Let $E/\mathbb{Q}$ be an elliptic curve defined over $\mathbb{Q}$. We define the following 6-tuple as Birch and Swinnerton-Dyer (BSD) data following $L$-functions and Modular Forms Database (LMFDB) [7]:
$\mathrm{BSD}(E/\mathbb{Q})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bigl{(}L(E,s),\,E(\mathbb{Q}),\,\mathrm{Reg}(E/\mathbb{Q}),\,\Omega_{E},\,c_{E/\mathbb{Q}_{p}},\,\Sha(E/\mathbb{Q})\bigr{)}.$
Here, the definition of each component is as follows.
  • $L(E,s)$: The $L$-function of $E/\mathbb{Q}$.
  • $E(\mathbb{Q})$: Mordell–Weil group of $E/\mathbb{Q}$, which is isomorphic to $\mathbb{Z}^{\text{rank}(E/\mathbb{Q})}\times E(\mathbb{Q})_{\text{tor}}$.
  • $\text{Reg}(E/\mathbb{Q})$: The regulator of $E/\mathbb{Q}$.
  • $\Omega_{E}$: The real period of $E$, given by $\Omega_{E}=\int_{E(\mathbb{R})}|\omega|$ where $\omega$ is the invariant differential of global minimal model of $E/\mathbb{Q}$.
  • $c_{E/\mathbb{Q}_{p}}$: The Tamagawa number at a prime number $p$.
  • $\Sha(E/\mathbb{Q})$: The Tate-Shafarevich group of $E/\mathbb{Q}$, conjecturally finite.
When $E_{1}\cong E_{2}$ over $\mathbb{Q}$, we have $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$.
Question.
When $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$, does $E_{1}\cong E_{2}$ over $\mathbb{Q}$ hold?
Having the same $L$-function is equivalent to having the same $\ell$-adic Galois representation for some prime $\ell$ and to being isogenous over $\mathbb{Q}$. Thus, we should choose $(E_{1},E_{2})$ to be from the same isogeny class over $\mathbb{Q}$.
It is known that the quantity
$\frac{\Omega_{E}\cdot\text{Reg}(E/\mathbb{Q})\cdot\#\Sha(E/\mathbb{Q})\cdot\prod_{p}c_{E/\mathbb{Q}_{p}}}{\#E(\mathbb{Q})_{\text{tor}}^{2}}$
remains invariant under isogeny under the assumption that $\Sha(E/\mathbb{Q})$ is finite ([12], Theorem 2.1), though none of individual terms in the product need be the same for isogenous curves.
In the high-dimensional case, Jamie Bell proved the existence of pairs of abelian varieties of dimension $p-1$ (where the class number of $\mathbb{Q}(\zeta_{p})$ is not $1$, so $p\geq 23$), defined over $\mathbb{Q}$, such that over every number field they share the same BSD invariants (except for the real period), the same $n$-Selmer groups for all $n$, and the same Tate modules [2]. The pair becomes isomorphic over $\overline{\mathbb{Q}}$. There are many coprime degree isogenies between this pair of abelian varieties. However, for two non-isomorphic elliptic curves over $\mathbb{Q}$, there cannot exist two or more isogenies of coprime degrees. Therefore, we need to examine the $q$-primary component of the invariants in the case of elliptic curves connected by prime degree $q$ isogenies.
Additionally, recent research has been conducted on classifying pairs of isogenous elliptic curves with the same minimal discriminant, called discriminant twins [1]. Our research considers whether elliptic curves can be identified by the coincidence of even more invariants. We prove the following theorem.
Theorem 1.1.
There exist infinitely many pairs $(E_{1},E_{2})$ of non-isomorphic elliptic curves over $\mathbb{Q}$ such that $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$, and the Kodaira symbols at every prime and the minimal discriminants are the same.
As noted in Remark 5.8, it has been proved that—up to twists—there is exactly one pair of 2‑isogenous elliptic curves that share the same BSD invariants, Kodaira symbols, and minimal discriminants. In summary, up to quadratic twist, there exists exactly one 2‑isogenous pair of elliptic curves over $\mathbf{Q}$ that cannot be distinguished by their BSD invariants, Kodaira symbols, minimal discriminants. And such indistinguishable pairs occur with positive density among quadratic twists of elliptic curves satisfying the sufficient condition of Proposition 5.4.
1.1. Acknowledgements
I would like to express my deepest gratitude to Prof. Nobuo Tsuzuki for his many constructive discussions and continual encouragement throughout this work. I also thank Wataru Kai and Yoshinori Mishiba, as well as the other participants of the Tohoku University Number Theory Seminar, for their valuable comments.
2. Notation and outline of the proof
2.1. Notation
  • For an abelian group $A$ and a prime number $p$,
    $A[p]$ :=$:=\{\,a\in A\mid pa=0\}$ ($p$‑torsion subgroup)
    $A[p^{\infty}]$ :=$:=\{\,a\in A\mid\exists\,n\geq 1,\;p^{n}a=0\}$ ($p$‑primary component)
  • $\Delta_{E}$: minimal discriminant of elliptic curve $E$.
  • $j(E)$: $j$-invariant of elliptic curve $E$.
  • $E^{D}/\mathbb{Q}$: For the Weierstrass model $E/\mathbb{Q}:y^{2}=x^{3}+ax^{2}+bx+c$, the quadratic twist of $E/\mathbb{Q}$ by a square-free integer $D$ is an elliptic curve defined by the equation
    $E^{D}:y^{2}=x^{3}+aDx^{2}+bD^{2}x+cD^{3}.$
  • For an isogeny $\phi$ between abelian varieties, $\hat{\phi}$ denotes the dual isogeny of $\phi$.
  • For an elliptic curve $E/\mathbb{Q}$, another elliptic curve $E^{\prime}/\mathbb{Q}$ and a nonzero isogeny $\phi:E\to E^{\prime}$, $\mathrm{Sel}^{\phi}(E/\mathbb{Q})$ is the $\phi$-Selmer group. When $E=E^{\prime}$ and $\phi=[n]$ (multiplication-by-$n\geq 1$ map), we denote its Selmer group by $\mathrm{Sel}^{n}(E/\mathbb{Q})$.
  • $\mathrm{rank}_{an}(E/\mathbb{Q})\coloneqq\mathrm{ord}_{s=1}L(E,s)$: analytic rank of $E/\mathbb{Q}$.
  • $\mathbb{Q}(E[2])$: the minimal extension of $\mathbb{Q}$ containing every point in $E(\bar{\mathbb{Q}})[2]$.
  • $v_{p}(\cdot)$: For a prime number $p$, $v_{p}$ denotes the $p$-adic valuation.
2.2. Outline of the proof
First, Proposition 5.4 provides a sufficient condition for a non-isomorphic, 2-isogenous pair of elliptic curves to share all six BSD invariants.
The LMFDB already contains several pairs satisfying this condition, and one can easily confirm that these pairs indeed have the same BSD invariants and Kodaira symbols by Proposition 5.4. Applying quadratic twists then yield infinitely many further pairs; the existence of suitable twists follows in general from the Smith density theorem or classical two descent (Proposition 6.1) in the special case that we have sought.
To align the Tate–Shafarevich groups, we simultaneously trivialize their 2-torsion subgroups. By Proposition 5.3, the entire Tate–Shafarevich groups therefore coincide.
The logical flow of each section is roughly as follows. The parts enclosed in $\square$ are the main logic flow.
Proposition 5.3Proposition 5.4Theorem 5.6Example 5.5Section 4Theorem 1.1 Proposition 3.4 or Proposition 6.1 Theorem 3.2
3. Simultaneous trivialization of $2$-torsion subgroups of the Tate-Shafarevich groups of two elliptic curves related via a degree $2$ isogeny
Definition 3.1.
A degree $2$ isogeny $\phi:E_{1}\to E_{2}$ defined over $\mathbb{Q}$ is called balanced if it satisfies $\mathbb{Q}(E_{1}[2])=\mathbb{Q}(E_{2}[2])$.
Recently, Smith’s result [11] has enabled us to determine the density of the joint distribution of $\phi$-Selmer groups and $\hat{\phi}$-Selmer groups when $\phi$ is balanced. Although Proposition 6.1 alone would already suffice to prove Theorem 1.1, we prove the more general Proposition 3.4, of which Proposition 6.1 is a special case. Combined with Smith’s result, we can obtain a positive-density family of desired quadratic twists from any pair of elliptic curves satisfying conditions $(1)$ and $(3)$ in Proposition 5.4.
Theorem 3.2 (Theorem 1.10, [11]).
Let $\phi:E_{1}\to E_{2}$ be a degree two balanced isogeny defined over $\mathbb{Q}$. Choose a nonzero integer $D_{0}$. Suppose that square-free integer $D$ satisfies
$DD_{0}\in(\mathbb{Q}_{v}^{\times})^{2}\ \text{for every}\ v\in S\ \text{where}\ S\coloneqq\{v:\text{place of}\ \mathbb{Q}\mid v\mid 2\Delta_{E_{1}}\Delta_{E_{2}}\infty\}.$
Take $u=\dim_{\mathbb{F}_{2}}\operatorname{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})-\dim_{\mathbb{F}_{2}}\operatorname{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})$. Let
$X_{E}(D_{0},H)\;:=\;\Bigl{\{}D\in\mathbb{Z}:\text{square-free}\;\Bigm{|}\;|D|\leq H,\;D\,D_{0}\in({\mathbb{Q}_{v}}^{\times})^{2}\text{ for every }v\in S\Bigr{\}}.$
Then, for any $r\geq 0$,
$\lim_{H\to\infty}\frac{\#\{D\in X_{E}(D_{0},H)\,:\,\,\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{D})=r+1\}}{\#X_{E}(D_{0},H)}=\alpha_{r,u}.$
where $\alpha_{r,u}$ denotes the limit, as $n\to\infty$, of the probability that a uniformly random $(n-u)\times n$ matrix over $\mathbb{F}_{2}$ has rank exactly $n-r$.
Remark 3.3.
For any square-free $D$ satisfying $DD_{0}\in(\mathbb{Q}_{v}^{\times})^{2}$ for every $v\in S$, $u$ is independent of the choice of $D$. Define the Tamagawa ratio by
$\tau^{D}\coloneqq\frac{\#\,\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})}{\#\,\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})}.$
Then
$\tau^{D}=\prod_{v}\left(\frac{1}{2}\,\#\,W_{v}(\phi,D)\right),$
where $W_{v}(\phi,D)\;:=\;\ker\mskip-3.0mu\Bigl{(}H^{1}(\mathbb{Q}_{v},E_{1}^{D}[\phi])\longrightarrow H^{1}(\mathbb{Q}_{v},E_{1}^{D})\Bigr{)}$  (see Theorem 3.5 of [6]).
When $E^{D}/\mathbb{Q}$ has good reduction at an odd prime $v$, $W_{v}(\phi,D)\cong H^{1}_{\mathrm{nr}}(\mathbb{Q}_{v},E^{D}[\phi])\cong H^{1}(\widehat{\mathbb{Z}},\mathbb{Z}/2\mathbb{Z})\cong\mathbb{Z}/2\mathbb{Z}$ where $H^{1}_{\mathrm{nr}}(\mathbb{Q}_{v},E^{D}[\phi])$ is the unramified cohomology. When $E/\mathbb{Q}$ has good reduction at an odd prime $v$ and $v(D)=1$, $W_{v}(\phi,D)\cong\mathbb{Z}/2\mathbb{Z}$ (see section 3.1 of [11]). The remaining cases are where $E/\mathbb{Q}$ has bad reduction at $v$ or $v=2$ or $v=\infty$. Since $DD_{0}\in(\mathbb{Q}_{v}^{\times})^{2}$ for every $v\in S$, $E_{1}^{D}[\phi]\cong E_{1}^{D_{0}}[\phi]$ and $W_{v}(\phi,D)\cong W_{v}(\phi,D_{0})$. Summarizing the above, $\tau^{D}=\tau^{D_{0}}$. We use this remark to simultaneously control $\Sha(E_{1}^{D}/\mathbb{Q})$ and $\Sha(E_{2}^{D}/\mathbb{Q})$ in Section 6.
Proposition 3.4.
Suppose $E_{i}/\mathbb{Q}\ (i\in\{1,2\})$ is an elliptic curve with $E_{i}(\mathbb{Q})[2]\simeq\mathbb{Z}/2\mathbb{Z}$. Suppose that $E_{1}$ and $E_{2}$ are isogenous via a degree $2$ balanced isogeny $\phi$. Suppose that $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}/\mathbb{Q})$. Then, there exist infinitely many square-free integers $D$ such that
  • $\Sha(E^{D}_{1}/\mathbb{Q})[2]=\Sha(E^{D}_{2}/\mathbb{Q})[2]=0$
  • $\text{gcd}(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$
  • $\mathrm{rank}_{an}(E_{1}^{D}/\mathbb{Q})\equiv 0\bmod 2$.
Proof.
Let us take the integral model $E_{1}:y^{2}=x^{3}+ax^{2}+bx,E_{2}:y^{2}=x^{3}-2ax^{2}+(a^{2}-4b)x\ (a,b\in\mathbb{Z})$ and let $\phi:E_{1}\to E_{2},\phi(x,y)=(\frac{y^{2}}{x^{2}},\frac{y(b-x^{2})}{x^{2}})$. For $D\in\mathbb{Z}$, let $e_{D}=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E^{D}_{1}/\mathbb{Q})$ and $f_{D}=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E^{D}_{2}/\mathbb{Q})$. There exists an exact sequence:
$0\to E^{D}_{2}(\mathbb{Q})[\hat{\phi}]/\phi(E_{1}^{D}(\mathbb{Q})[2])\to\mathrm{Sel}^{\phi}(E^{D}_{1}/\mathbb{Q})\to\mathrm{Sel}^{2}(E^{D}_{1}/\mathbb{Q})\stackrel{{\scriptstyle\phi}}{{\to}}\mathrm{Sel}^{\hat{\phi}}(E^{D}_{2}/\mathbb{Q})$
(see [[9], lemma 9.1]). Thus,
$\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{2}(E_{1}^{D}/\mathbb{Q})\leq e_{D}+f_{D}-\dim_{\mathbb{F}_{2}}\dfrac{E^{D}_{2}(\mathbb{Q})[\hat{\phi}]}{\phi(E^{D}_{1}(\mathbb{Q})[2])}.$
By Theorem 3.2, there exist infinitely many square-free integers $D$ such that $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$ and $e_{D}=f_{D}=1$ by taking $D_{0}=1$ and $r=0$. Indeed, with $D_{0}=1$, we have $u=e_{D}-f_{D}=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})=0$. Also, $D\in(\mathbb{Q}_{v}^{\times})^{2}$ for every $p\mid 2\Delta_{E_{1}}\Delta_{E_{2}}$ implies $\text{gcd}(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$. Hence, it remains only to show that $\alpha_{0,0}>0$. In fact,
$\alpha_{0,0}=\prod_{s=1}^{\infty}(1-2^{-s})=0.288\ldots>0$
as required.
For this square-free integer $D$, we have
(3.5) #$\#\Sha(E^{D}_{i}/\mathbb{Q})[2]=\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})-\dim_{\mathbb{F}_{2}}E^{D}_{i}(\mathbb{Q})/2E^{D}_{i}(\mathbb{Q})\leq 1-1=0$
for $i\in\{1,2\}$. Additionally, since $\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}\leq\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})\leq 2$ and #$\#\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}=2^{\mathrm{rank}(E^{D}_{i}/\mathbb{Q})+1}$, we have $\mathrm{rank}(E^{D}_{i}/\mathbb{Q})=0$ for $i\in\{1,2\}$.
Note that the condition $\mathrm{rank}_{an}(E_{1}^{D}/\mathbb{Q})=\mathrm{rank}_{an}(E_{2}^{D}/\mathbb{Q})\equiv 0\bmod 2$ follows automatically from $\Sha(E^{D}_{1}/\mathbb{Q})[2]\cong\Sha(E^{D}_{2}/\mathbb{Q})[2]=0$ and $\mathrm{rank}(E^{D}_{1}/\mathbb{Q})=\mathrm{rank}(E^{D}_{2}/\mathbb{Q})=0$. Indeed, let $\omega_{E}$ be a root number of $E/\mathbb{Q}$. The condition that $\mathrm{rank}(E^{D}_{1}/\mathbb{Q})=0$ and $\Sha(E^{D}_{1}/\mathbb{Q})[2]=0$ tell us that $2^{\infty}$-Selmer rank of $E^{D}_{1}/\mathbb{Q}$, that is $\mathrm{rank}(E^{D}_{1}/\mathbb{Q})+\text{number of copies of }\mathbb{Q}_{2}/\mathbb{Z}_{2}\text{ in }\Sha(E^{D}_{1}/\mathbb{Q})$, is zero. By [8], Theorem 1.4, this implies $\omega(E^{D}_{1}/\mathbb{Q})=1$, thus $\mathrm{rank}_{an}(E^{D}_{1}/\mathbb{Q})$ is even. ∎
4. Kodaira symbol, Tamagawa number and real period
In this chapter, we recall discriminant twins and results on Kodaira symbols , Tamagawa numbers, and real periods.
Theorem 4.1.
Let $E_{1}$ and $E_{2}$ be isogenous elliptic curves over $\mathbb{Q}$, and assume $E_{1}$ has good reduction at $2$. The following are equivalent:
  1. (1)
    $E_{1}$ and $E_{2}$ have the same Kodaira symbol at every prime $\ell$;
  2. (2)
    $\Delta_{E_{1}}=\pm\Delta_{E_{2}}$.
Proof.
($(1)\Rightarrow(2)$) Suppose all Kodaira symbols are the same. Ogg’s formula states that
$v_{l}(\Delta_{E})=f_{l}+m_{l}-1,$
where $f_{l}$ is the conductor exponent and $m_{l}=\#\tilde{\epsilon}(\overline{\mathbb{F}_{l}})/\#\tilde{\epsilon}_{0}(\overline{\mathbb{F}_{l}})$. Here, $\tilde{\epsilon}$ is the special fiber of the Néron minimal model of $E/\mathbb{Q}$, $\tilde{\epsilon}_{0}$ is its connected component of the identity, and $\Delta_{E}$ is the minimal discriminant of $E/\mathbb{Q}$. Since $E_{1}$ and $E_{2}$ are isogenous, the conductor exponent $f_{l}$ is an isogeny invariant, so $f_{l,E_{1}}=f_{l,E_{2}}$ for all primes $l$. Since $E_{1}$ and $E_{2}$ have the same Kodaira symbols at all primes $l$, we have $m_{l,E_{1}}=m_{l,E_{2}}$ for all primes $l$. Therefore, $v_{l}(\Delta_{E_{1}})=v_{l}(\Delta_{E_{2}})$ for all primes $l$. This implies $\Delta_{E_{1}}=\pm\Delta_{E_{2}}$.
($(1)\Leftarrow(2)$) Suppose $\Delta_{E_{1}}=\pm\Delta_{E_{2}}$. Then for all primes $\ell$, we have $v_{\ell}(\Delta_{E_{1}})=v_{\ell}(\Delta_{E_{2}})$. By Ogg’s formula, this means $f_{\ell,E_{1}}+m_{\ell,E_{1}}-1=f_{\ell,E_{2}}+m_{\ell,E_{2}}-1$. Since $E_{1}$ and $E_{2}$ are isogenous, $f_{\ell,E_{1}}=f_{\ell,E_{2}}$, which implies $m_{\ell,E_{1}}=m_{\ell,E_{2}}$. By lemma 2.16 of [1], $E_{1}$ has potentially good reduction at $l>2$. In the additive potentially good case, the Kodaira symbol is uniquely determined by $m_{\ell}$ (See [10], IV.9, Table 4.1). Since both invariants are equal, the Kodaira symbols are the same. ∎
Theorem 4.2.
Suppose $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ are elliptic curves such that there exists a degree $2$ isogeny between them. Let $l\geq 3$ be a prime and $F$ be a quadratic extension of $\mathbb{Q}_{l}$ over which $E_{1}$ (or $E_{2}$) has good reduction. If $\frac{\Delta_{E_{1}}}{\Delta_{E_{2}}}\in N_{F/\mathbb{Q}_{l}}$, then $c_{E_{1}/\mathbb{Q}_{l}}=c_{E_{2}/\mathbb{Q}_{l}}$.
Proof.
See [3], Table 1 and Theorem 6.1. ∎
Proposition 4.3.
Let $p$ be a prime number. Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous by a degree $p$ isogeny over $\mathbb{Q}$. Assume that $E_{1}$ and $E_{2}$ have the same Tamagawa numbers at all primes and $\mathrm{rank}_{an}(E_{1}/\mathbb{Q})\equiv 0\bmod 2$. Then, $\Omega_{E_{1}}=\Omega_{E_{2}}$.
In particular, if $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$ and $E_{1}$ and $E_{2}$ are isogenous by degree $p$ isogeny, then $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ have even analytic rank.
Proof.
By Theorem 8.2 of [3],
$\Omega_{E_{1}}=\Omega_{E_{2}}\Leftrightarrow\sum_{l}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{l}}}{c_{E_{2}/\mathbb{Q}_{l}}})\equiv\text{ord}_{s=1}L(E_{1},s)\bmod 2.$
Since Tamagawa numbers at all primes are the same, $\sum_{l}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{l}}}{c_{E_{2}/\mathbb{Q}_{l}}})=0$ holds. Thus, the equality $\Omega_{E_{1}}=\Omega_{E_{2}}$ is equivalent to the congruence $\text{ord}_{s=1}L(E_{1},s)\equiv 0\bmod 2$, which is true.
5. Infinitely many pairs of elliptic curves sharing the same BSD data
Theorem 5.1.
If $E$ is an elliptic curve over $\mathbb{Q}$, then there are only finite number of square-free integers $D$ such that $E_{D}$ has a torsion point of order $>2$.
Proof.
See Proposition 1 of [4]. ∎
Lemma 5.2.
Let $F$ be a covariant functor from the category of abelian varieties to the category of abelian groups such that $F([n])=[n]$ for all positive integer $n$ where $[n]$ is the multiplication by $n$ map. Let $p$ be a prime, and suppose $E_{1}$ and $E_{2}$ are connected by an isogeny $\phi$ of degree $p$. Assume that for any elliptic curve $E$, $F(E)$ is a torsion group. If $F(E_{1})[p^{\infty}]\cong F(E_{2})[p^{\infty}]$ holds, then $F(E_{1})\cong F(E_{2})$ holds.
Proof.
Since $\phi\circ\hat{\phi}=[p]$ and $\hat{\phi}\circ\phi=[p]$, we have $F(\phi)\circ F(\hat{\phi})=[p]$ and $F(\hat{\phi})\circ F(\phi)=[p]$. For a prime $l\neq p$ and for all $m$, since $\gcd(p,l)=1$, $F([p])=p:F(E_{1})[l^{\infty}]\to F(E_{2})[l^{\infty}]$ is bijective, it follows that $F(\phi):F(E_{1})[l^{\infty}]\to F(E_{2})[l^{\infty}]$ is an isomorphism. Since $F(E_{1})$ and $F(E_{2})$ are torsion groups, we have
$F(E_{1})\cong\bigoplus_{q:\text{prime}}F(E_{1})[q^{\infty}]\cong\bigoplus_{q:\text{prime}}F(E_{2})[q^{\infty}]\cong F(E_{2}).$
Proposition 5.3.
Let $p$ be a prime. Let $E_{1}/\mathbb{Q}$, $E_{2}/\mathbb{Q}$ be elliptic curves over $\mathbb{Q}$ that are isogenous of degree $p$.
(1) Suppose $E_{1}(\mathbb{Q})[p^{\infty}]\cong E_{2}(\mathbb{Q})[p^{\infty}]$ holds. Then $E_{1}(\mathbb{Q})\cong E_{2}(\mathbb{Q})$.
(2) Let $p$ be a prime and suppose $\Sha(E_{1}/\mathbb{Q})[p^{\infty}]\cong\Sha(E_{2}/\mathbb{Q})[p^{\infty}]$ holds. Then, without assuming the finiteness of the Tate–Shafarevich group, $\Sha(E_{1}/\mathbb{Q})\cong\Sha(E_{2}/\mathbb{Q})$ holds.
Proof.
(1) The rank is invariant under isogeny. By Lemma 5.2, the torsion parts are also isomorphic. By the Mordell–Weil theorem, $E_{1}(\mathbb{Q})\cong E_{2}(\mathbb{Q})$. (2) Any element $[C]$ of the Tate–Shafarevich group vanishes when lifted to some degree $n$ extension, we have $n\cdot[C]=0$. Therefore, the Tate-Shafarevich group is a torsion group. By Lemma 5.2, we have $\Sha(E_{1}/\mathbb{Q})\cong\Sha(E_{2}/\mathbb{Q})$. ∎
Proposition 5.4.
Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous via a degree–$2$ isogeny defined over $\mathbb{Q}$. Assume that
  1. (1)
    $E_{1}(\mathbb{Q})[2^{\infty}]\;\cong\;E_{2}(\mathbb{Q})[2^{\infty}]$.
  2. (2)
    $\Sha(E_{1}/\mathbb{Q})[2^{\infty}]\;\cong\;\Sha(E_{2}/\mathbb{Q})[2^{\infty}]$,
    $\operatorname{rank}_{\mathrm{an}}\mskip-3.0muE_{1}(\mathbb{Q})\equiv 0\pmod{2}$, and $\operatorname{rank}E_{1}(\mathbb{Q})=0$.
  3. (3)
    $E_{1}$ has good reduction at $2$ and $\Delta_{E_{1}}=\Delta_{E_{2}}$.
Then $\mathrm{BSD}(E_{1}/\mathbb{Q})=\mathrm{BSD}(E_{2}/\mathbb{Q})$, and $E_{1}$ and $E_{2}$ have the same Kodaira symbols at every prime.
Proof.
Since $E_{1}$ and $E_{2}$ are isogenous over $\mathbb{Q}$, $L(E_{1},s)=L(E_{2},s)$.
By Theorem 4.1, the Kodaira symbols are the same. By Theorem 4.2, Tamagawa numbers at $l\geq 3$ are the same because $\Delta_{E_{1}}=\Delta_{E_{2}}$. For the period, $\Omega_{E_{1}}=\Omega_{E_{2}}$ follows from Proposition 4.3. Because $\mathrm{rank}(E_{i}/\mathbb{Q})=0$, $\mathrm{Reg}(E_{i}/\mathbb{Q})=1$ for $i\in\{1,2\}$.
By Proposition 5.3, we have $E_{1}(\mathbb{Q})\cong E_{2}(\mathbb{Q})$ and $\Sha(E_{1}/\mathbb{Q})\cong\Sha(E_{2}/\mathbb{Q})$.
Example 5.5.
Examples of pairs $(E_{a},E_{b})$ of elliptic curves that satisfy conditions (1), (2), and (3) of Proposition 5.4 are $(38025.ck1,38025.ck2)$ (this is the same as the pair described in Proposition 6.1), $(38025.t1,38025.t2)$, $(38025.u1,38025.u2)$, $(38025.cl1,38025.cl2)$, $(207025.bf1,207025.bf2)$, $(207025.bu1,207025.bu2)$, and $(207025.cj1,207025.cj2)$. As explained in Remark 5.8, these are unique pairs of 2-isogenous elliptic curves that share the same minimal discriminants, up to twists.
Theorem 5.6.
Suppose that $E_{1}(\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}$ and $E_{1}/\mathbb{Q}$ has good reduction at $2$. Suppose $(E_{1},E_{2})$ is a pair of a balanced isogenous elliptic curves with $E_{1}(\mathbb{Q})[2^{\infty}]\cong E_{2}(\mathbb{Q})[2^{\infty}]$, and $\Delta_{E_{1}}=\Delta_{E_{2}}$. Additionally, suppose $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}/\mathbb{Q})$. Then there exist infinitely many square-free integers $D$ such that $\text{BSD}(E^{D}_{1}/\mathbb{Q})=\text{BSD}(E^{D}_{2}/\mathbb{Q})$ and Kodaira symbols at every prime and minimal discriminants are the same.
Proof.
For square-free integers $D$, $E^{D}_{1}$ and $E^{D}_{2}$ are isogenous over $\mathbb{Q}$. We prove that infinitely many square-free $D$ satisfy the following conditions .
  1. (1)
    $E^{D}_{1}(\mathbb{Q})[2^{\infty}]\cong E^{D}_{2}(\mathbb{Q})[2^{\infty}]$.
  2. (2)
    $\Sha(E^{D}_{1}/\mathbb{Q})[2^{\infty}]\cong\Sha(E^{D}_{2}/\mathbb{Q})[2^{\infty}]=0$,
    $\operatorname{rank}_{an}E^{D}_{1}(\mathbb{Q})\equiv 0\pmod{2}$ and $\mathrm{rank}(E^{D}_{1}/\mathbb{Q})=0$.
  3. (3)
    $\operatorname{ord}_{2}(\Delta_{E^{D}_{1}})=0$ and $\Delta_{E^{D}_{1}}=\Delta_{E^{D}_{2}}$.
Then, by Proposition 5.4, we are done.
For condition (1), For a square-free integer $D$, $E_{1}^{D}(\mathbb{Q})[2]\cong E_{2}^{D}(\mathbb{Q})[2]$. By Theorem 5.1, there are finitely many $D$ such that $E^{D}_{1}(\mathbb{Q})_{\text{tor}}$ or $E^{D}_{2}(\mathbb{Q})_{\text{tor}}$ acquire points of order greater than 4. Thus, for almost all square-free integers $D$, we have $E^{D}_{1}(\mathbb{Q})[2^{\infty}]\cong E^{D}_{2}(\mathbb{Q})[2^{\infty}]$. By Proposition 5.3 (2), $E^{D}_{1}(\mathbb{Q})\cong E^{D}_{2}(\mathbb{Q})$.
For (2), $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E^{D}_{1}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E^{D}_{2}/\mathbb{Q})$ by Remark 3.3. By Proposition 3.4, infinitely many square-free integer $D$ satisfy $\Sha(E^{D}_{1}/\mathbb{Q})[2]=\Sha(E^{D}_{2}/\mathbb{Q})[2]=0$, $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$ and $\mathrm{rank}_{an}(E^{D}_{1}/\mathbb{Q})\equiv 0\bmod 2$. For such $D$, by Proposition 5.3 (2), $\Sha(E^{D}_{1}/\mathbb{Q})\cong\Sha(E^{D}_{2}/\mathbb{Q})$.
For (3), let us take $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$. Note that $E^{D}_{1}/\mathbb{Q},E^{D}_{2}/\mathbb{Q}$ is global minimal when the model of $E_{1}/\mathbb{Q},E_{2}/\mathbb{Q}$ is global minimal because $(D,\Delta_{E_{1}}\Delta_{E_{2}})=1$. Also, $E^{D}_{1}$ has good reduction at $2$. Since $\frac{\Delta_{E^{D}_{2}}}{\Delta_{E^{D}_{1}}}=\frac{\Delta_{E_{2}}}{\Delta_{E_{1}}}=1$, we have $\Delta_{E^{D}_{1}}=\Delta_{E^{D}_{2}}$.
Theorem 5.7.
There exist infinitely many pairs $(E_{1},E_{2})$ of non-isomorphic elliptic curves over $\mathbb{Q}$ such that $j(E_{1})\neq j(E_{2})$, $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$, and the Kodaira symbols at every prime and the minimal discriminant is the same.
Proof.
The pairs in Example 5.5 satisfy the hypothesis of Proposition 5.4. By Theorem 5.6, infinitely many pairs $(E^{D}_{a},E^{D}_{b})$ are the pairs we have sought. ∎
Finally, we see that these examples are very rare.
Remark 5.8.
In [1], pairs of isogenous elliptic curves over $\mathbb{Q}$ with the same minimal discriminants are called discriminant twins, and examples of these are classified in table 5 of [1]. Discriminant twins over $\mathbb{Q}$ are extremely rare. Recent work [1] proved that there is only one pair of discriminant twins among 2-isogenous curves up to twists: $(4225.h1,4225.h2)$. The curve 4225.h1 is a twist of our 38025.ck1 by $-195$. For distinct square–free integer $D_{1},D_{2}$, $E^{D_{1}},E^{D_{2}}$ do not have the same BSD invariants since the real period $\Omega_{E^{D_{1}}}=\frac{\Omega_{E}}{\sqrt{|D_{1}|}},\Omega_{E^{D_{2}}}=\frac{\Omega_{E}}{\sqrt{|D_{2}|}}$ never coincide.
6. Explicit infinite families of pairs
In this chapter, independently of Smith’s results, we give a concrete infinite family using classical 2-descent (see for example, [10], X, Remark 3.7 and Proposition 4.9) with a balanced isogeny case. The key point is that since this is a balanced isogeny case, we can simultaneously control $\Sha(E_{1}/\mathbb{Q})$ and $\Sha(E_{2}/\mathbb{Q})$ by using Remark 3.3.
6.1. Two descent for balanced isogenous pairs of elliptic curves
Proposition 6.1.
Let $E_{1}:y^{2}=x^{3}+25350x^{2}+2471625x$ and $E_{2}:y^{2}=x^{3}-50700x^{2}+632736000x$ (labeled as (38025.ck1, 38025.ck2) in LMFDB). Take a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$. For such $D$, $\Sha(E_{1}^{D}/\mathbb{Q})[2]=0$ and $\Sha(E_{2}^{D}/\mathbb{Q})[2]=0$.
Proof.
$E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ are isogenous over $\mathbb{Q}$ by $\phi:E_{1}\to E_{2}$, $(x,y)\mapsto\left(\frac{y^{2}}{x^{2}},\frac{y(2471625-x^{2})}{x^{2}}\right)$. Also, $\mathbb{Q}(E_{1}[2])=\mathbb{Q}(E_{2}[2])=\mathbb{Q}(\sqrt{65})$, so $\phi$ is balanced. Let us define
$S\coloneqq\{v\mid v\text{ divides }\deg\phi\}\cup\{v\mid v\text{ divides }\Delta_{E_{1}}\Delta_{E_{2}}\}\cup\{\infty\}=\{2,3,5,13,D,\infty\}$
and $\mathbb{Q}(S,2)\coloneqq\{\bar{d}\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}\mid v(d)\equiv 0\pmod{2}\text{ for all }v\notin S\}$. Then,
#$\#\mathbb{Q}(S,2)=\#\{(-1)^{n_{0}}2^{n_{1}}3^{n_{2}}5^{n_{3}}13^{n_{4}}D^{n_{5}}\mid 0\leq n_{1},n_{2},n_{3},n_{4},n_{5}\leq 1\}=64.$
Via the following isomorphism, the Selmer group can be calculated by checking the local solubility of torsors $C_{d}$.
$\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})\cong\left\{\bar{d}\in\mathbb{Q}(S,2)\;\middle|\;\begin{aligned} &\bar{C}_{d}(\mathbb{Q}_{v})\neq\emptyset,\forall v\in S,\text{where }\ \bar{C}_{d}\ \text{is the projective closure of}\\ &\text{affine curve}\ C_{d}:dy^{2}=d^{2}-50700dDx^{2}+632736000D^{2}x^{4}\ \text{in}\ \mathbb{P}^{3}\end{aligned}\right\}$
(see [10], X, Remark 3.7 and Proposition 4.9). Here, $50700=2^{2}\times 3\times 5^{2}\times 13^{2},632736000=2^{8}\times 3^{2}\times 5^{3}\times 13^{3}$.
$C_{d}$ has two points at infinity $\infty\coloneqq[0,0,\pm\sqrt{\frac{632736000}{d}},1]$ in $\mathbb{P}^{3}$ (see [10], X, Remark 3.7). Since $\sqrt{\frac{632736000}{65}}=3120$, we note that
(6.2) $65\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q}).$
For $64$ candidates $d\in\mathbb{Q}(S,2)$, we check which $d$ is in $\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$.
To prove our proposition, for $d\neq 1,65$, we find a place $p(d)\mid 2\cdot 3\cdot 5\cdot 13\cdot D\cdot\infty$ of $\mathbb{Q}$ for which $\bar{C}_{d}(\mathbb{Q}_{p(d)})=\emptyset$. The values of $d$ and $p(d)$ we choose are listed in the following table. As shown in the table below, there are no points at infinity. In what follows, we prove that $\bar{C}_{d}$ has no $\mathbb{Q}_{p(d)}$-rational points on the affine part.
$\bar{C}_{d}$ $p(d)$ third coordinate of points at $\infty$
$\bar{C}_{d}\,(d<0)$infty ±$\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{R}\ \text{since}\ d<0$
$\bar{C}_{d}\,(3\mid d)$ 3 ±$\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{Q}_{3}\ \text{since}\ v_{3}(\frac{632736000}{d})=2-n_{2}=1$
$\bar{C}_{2}$ 5 ±$\pm\sqrt{\frac{632736000}{2}}\notin\mathbb{Q}_{5}$
$\bar{C}_{13}$ 5 ±$\pm\sqrt{\frac{632736000}{13}}\notin\mathbb{Q}_{5}$
$\bar{C}_{D}$ 5 ±$\pm\sqrt{\frac{632736000}{D}}\notin\mathbb{Q}_{5}$
$\bar{C}_{5D}$ 13 ±$\pm\sqrt{\frac{632736000}{5D}}\notin\mathbb{Q}_{13}$
$\bar{C}_{2D}$ 13 ±$\pm\sqrt{\frac{632736000}{2D}}\notin\mathbb{Q}_{13}$
$\bar{C}_{10D}$ $D$$\pm\sqrt{\frac{63273600}{D}}\notin\mathbb{Q}_{D}$
$\bar{C}_{10}$ 2 ±$\pm\sqrt{\frac{632736000}{10}}\notin\mathbb{Q}_{2}$
When $d<0$, $\bar{C}_{d}\notin\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$. Thus, $n_{0}=0$. For $n_{2}\geq 1$, $v_{3}(dy^{2})=1+2v_{3}(y)$ is odd while $v_{3}(d^{2}-50700Ddx^{2}+632736000D^{2}x^{4})$ is even. Indeed, $v_{3}(d^{2})=2$, $v_{3}(50700Dd^{\prime}x^{2})=2+2v_{3}(x)$, $v_{3}(632736000D^{2}x^{4})=2+4v_{3}(x)$. When $v_{3}(x)$ is positive or negative, it is even. When $v_{3}(x)=0$, let $f(x)\coloneqq d^{2}-50700Ddx^{2}+632736000D^{2}x^{4}$ and $d=3d^{\prime}$. Now, $\frac{f(x)}{9}\in\mathbb{Z}_{3}$ and $\frac{f(x)}{9}=d^{\prime 2}-16900Dd^{\prime}x^{2}+70304000D^{2}x^{4}\equiv d^{\prime 2}+2Dd^{\prime}x^{2}+2D^{2}x^{2}\equiv 1+2Dd^{\prime}+2D^{2}\not\equiv 0\bmod 3$, so $v_{3}(f(x))=2$. Thus, $n_{2}$ should be 00. So, the candidate of $d\in\mathbb{Q}(S,2)$ such that $d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$ is of the form $2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$. Thus, $d=2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$. The left candidates of $d$ is $d=2,5,13,65,130,D,5D,13D,65D,10,26,10D,26D,130D$.
To check these candidates, it is sufficient to prove $d=2,13,D,5D,2D,10D,10$ does not belong to $\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$. Indeed,
$d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})\Leftrightarrow 65d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$
holds since $65\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$(see (6.2)).
By the following Lemma 6.3, we have $\text{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})=\{1,65\}\cong\mathbb{Z}/2\mathbb{Z}.$
2-descent shows that $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{17}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})=0$. In Remark 6.4, let us take $D_{0}=17$ so that $17D\in{\mathbb{Q}_{v}^{\times}}^{2}$. This condition is equivalent to $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $D>0$. Then, by Remark 6.4, since $\phi$ is balanced, we have $\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}$.
By (3.5) in the proof of Proposition 3.4, we have $\Sha(E^{D}_{1}/\mathbb{Q})[2]=\Sha(E^{D}_{2}/\mathbb{Q})[2]=0$ and $\mathrm{rank}(E_{1}^{D}/\mathbb{Q})=\mathrm{rank}(E_{2}^{D}/\mathbb{Q})=0$ for infinitely many square-free integer $D$.
Lemma 6.3.
Let $D>0$ be a prime number such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$. Then, for all $d\in\{2,13,D,5D,2D,10D,10\}$, there exists a place $p(d)\in\{v:\text{place of }\mathbb{Q}\mid v\text{ divides }2\cdot 3\cdot 5\cdot 13\cdot D\text{ or }v=\infty\}$ of $\mathbb{Q}$ such that the affine curve $C_{d}:dy^{2}=d^{2}-50700dDx^{2}+632736000D^{2}x^{4}$ does not have $\mathbb{Q}_{p(d)}$-rational points.
Proof.
For each $d$, take $p(d)$ and assume $C_{d}(\mathbb{Q}_{p(d)})\neq\emptyset$ to prove $C_{d}(\mathbb{Q}_{p(d)})=\emptyset$ by contradiction. $C_{d}$ is isomorphic to $y^{2}=d-50700Dx^{2}+\frac{632736000D^{2}}{d}x^{4}$. Let $f_{d}(x)\coloneqq d-50700Dx^{2}+\frac{632736000D^{2}}{d}x^{4}$. When $v_{p(d)}(f(x))$ is odd, we have $C_{d}(\mathbb{Q}_{p(d)})=\emptyset$ since $v_{p(d)}(y^{2})=2v_{p(d)}(y)$ is even.
For $d=2$, take $p(2)=5$. we have $v_{5}(-50700Dx^{2})=2+2v_{5}(x)$ and $v_{5}(316368000D^{2}x^{4})=3+4v_{5}(x)$. When $v_{5}(x)<0$, $v_{5}(f_{2}(x))=\text{min}\{0,2+2v_{5}(x),3+4v_{5}(x)\}=3+4v_{5}(x)$ is odd. When $v_{5}(x)\geq 0$, $y^{2}\equiv 2\pmod{5}$ cannot occur.
For $d=13$, take $p(13)=5$. $C_{13}:y^{2}=13-50700Dx^{2}+48672000D^{2}x^{4}$. Thus $v_{5}(x)\geq 0$ and $v_{5}(y)\geq 0$. By looking $C_{13}$ mod 5, we obtain $y^{2}\equiv 3\bmod 5$ cannot occur. If $v_{5}(x)<0$, since $v_{5}(f_{13}(x))=\text{min}\{0,2+2v_{5}(x),3+4v_{5}(x)\}=3+4v_{5}(x)$ is odd, $C_{13}(\mathbb{Q}_{5})=\emptyset$.
For $d=D$, take $p(D)=5$. $C_{D}:y^{2}=D-50700Dx^{2}+632736000Dx^{4}$. Since $v_{5}(D)=0,v_{5}(-50700Dx^{2})=2+2v_{5}(x),v_{5}(632736000Dx^{4})=3+4v_{5}(x)$. When $v_{5}(x)\geq 0$, $y^{2}\equiv D\bmod 5$. This contradicts that $\left(\frac{D}{5}\right)=-1$. When $v_{5}(x)<0$, $v_{5}(f_{D}(x))=\text{min}\{0,2+2v_{5}(x),3+4v_{5}(x)\}=3+4v_{5}(x)$ is odd and thus $C_{D}(\mathbb{Q}_{5})=\emptyset$.
[…]
For $d=2D$, take $p(2D)=13$. $C_{2D}:y^{2}=2D-50700Dx^{2}+316368000Dx^{4}$. When $v_{13}(x)\geq 0$, by looking $C_{2D}\bmod 13$, $y^{2}\equiv 2D\bmod 13$, which cannot occur. Since $v_{13}(x)<0$, $v_{13}(f_{2D}(x))=\text{min}\{0,2+2v_{13}(x),3+4v_{13}(x)\}=3+4v_{13}(x)$ is odd, $C_{2D}(\mathbb{Q}_{13})=\emptyset$.
For $d=10D$, take $p(10D)=D$. $C_{10D}:y^{2}=10D-50700Dx^{2}+63273600Dx^{4}$. When $v_{D}(x)=0$, we claim $v_{D}(g(x))=0$ where $g(x)\coloneqq 10-50700x^{2}+63273600x^{4}$. The discriminant $\Delta$ of the polynomial $1063273600t^{2}-50700t+$ 10 is $\Delta=2^{4}\times 3^{2}\times 5^{3}\times 13^{3}$. Since $\left(\frac{\Delta}{D}\right)=\left(\frac{5\times 13}{D}\right)=\left(\frac{5}{D}\right)\left(\frac{13}{D}\right)=-1$, $v_{D}(g(x))=0$ because $g(x)\equiv 0\bmod D$ does not have solutions. Thus, $v_{D}(Dg(x))=1$ is odd. When $v_{D}(x)$ is positive or negative, since $v_{D}(f_{10D}(x))$ is odd, we have $C_{10D}(\mathbb{Q}_{D})=\emptyset$.
For $d=10$, take $p(10)=2$. Then $C_{10}:y^{2}=10-50700Dx^{2}+63273600D^{2}x^{4}$ and we have $v_{2}(y^{2})=2v_{2}(y),v_{2}(50700Dx^{2})=2+2v_{2}(x)$, and $v_{2}(63273600D^{2}x^{4})=7+4v_{2}(x)$. When $v_{2}(x)=-1$, by substituting $Z=2x$, $C_{10}:y^{2}=10-12675DZ^{2}+3954600D^{2}Z^{4}$ and $v_{2}(Z)=0$. Since $y^{2}\equiv 2-3\times 1\times 1\equiv 7\bmod 8$ does not hold. When $v_{2}(x)=-2$, by substituting $W=4x$, $C_{10}:4y^{2}=40-12675DW^{2}+988650D^{2}W^{4}$. Since $4y^{2}\equiv 0-3\times 1\times 1+2\equiv 7\bmod 8$, $v_{2}(x)$ is not $-2$. The left case is $v_{2}(x)\leq-2$ or $v_{2}(x)\geq 0$. Since $v_{2}(10-12675DZ^{2}+3954600D^{2}Z^{4})$ is odd, we have $C_{10}(\mathbb{Q}_{2})=\emptyset$. ∎
6.2. Pairs of elliptic curves sharing the BSD invariants, Kodaira symbols and minimal discriminants
Combining Proposition 5.4 and Proposition 6.1, for a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$, and $D$ is large enough so that $E_{1}^{D}(\mathbb{Q})_{\text{tor}}$ and $E_{2}^{D}(\mathbb{Q})_{\text{tor}}$ do not have any elements of order $>2$, we have $\text{BSD}(E_{1}^{D}/\mathbb{Q})=\text{BSD}(E_{2}^{D}/\mathbb{Q})$ and the Kodaira symbols of $E_{1}$ and $E_{2}$ are the same. We provide a table of pairs of non-isomorphic elliptic curves that share the same BSD invariants.
Elliptic curve $y^{2}=x^{3}+25350Dx^{2}+2471625D^{2}x$ $y^{2}=x^{3}-50700Dx^{2}+632736000D^{2}x$
j-invariant $257^{3}$ $17^{3}$
Minimal discriminant $3^{6}\cdot 5^{9}\cdot 13^{9}\cdot D^{6}$ $3^{6}\cdot 5^{9}\cdot 13^{9}\cdot D^{6}$
Mordell–Weil group $\mathbb{Z}/2\mathbb{Z}$ $\mathbb{Z}/2\mathbb{Z}$
Regulator 1 1
Real period $\dfrac{0.209\cdots}{\sqrt{|D|}}$ $\dfrac{0.209\cdots}{\sqrt{|D|}}$
Tamagawa number $2(p=3),2(p=5),2(p=13)$ $2(p=3),2(p=5),2(p=13)$,
$\ddagger$(other prime $p$) $\ddagger$(other prime $p$)
Kodaira symbol $I_{0}^{*}(p=3),III^{*}(p=5),III^{*}(p=13)$, $I_{0}^{*}(p=3),III^{*}(p=5),III^{*}(p=13)$
$\ddagger$(other prime $p$) $\ddagger$(other prime $p$)
$\Sha(E/\mathbb{Q})[2^{\infty}]$ 0 0
$\Sha(E/\mathbb{Q})$dagger $\dagger$
For $\dagger$, note that $\Sha(E/\mathbb{Q})[l^{\infty}]$ remains unknown, but they agree as abelian groups.
Theshaded areas in the table indicate the conditions from Proposition 5.4 and this ensures all BSD invariants and the Kodaira symbol coincide.
For $\ddagger$, the Kodaira symbol at $p\mid D,p\nmid 2\cdot 3\cdot 5\cdot 13$ is $I_{0}^{*}$.
Remark 6.4.
Among these $D$, it would be an interesting problem to consider whether there exist infinitely many $D$ such that $\Sha(E_{1}^{D}/\mathbb{Q})\cong\Sha(E_{2}^{D}/\mathbb{Q})$ is non-trivial.

### v2

@@
 Abstract.
-We prove that there exist infinitely many pairs of non-isomorphic elliptic curves over $\mathbb{Q}$ sharing the same BSD invariants—including their $L$-functions, Mordell–Weil groups, Tate–Shafarevich groups, Tamagawa numbers, regulators, and real periods—and Kodaira symbols and minimal discriminants, while having distinct $j$-invariants.
+We prove that there exist infinitely many pairs of non-isomorphic elliptic curves over $\mathbb{Q}$ sharing the same BSD invariants—including their Mordell–Weil groups, Tate–Shafarevich groups, Tamagawa numbers, regulators, and real periods—and Kodaira symbols and minimal discriminants, while having distinct $j$-invariants.
 1. Introduction
-Let $E/\mathbb{Q}$ be an elliptic curve defined over $\mathbb{Q}$.
-We define the following 6-tuple as Birch and Swinnerton-Dyer (BSD) data following $L$-functions and Modular Forms Database (LMFDB) [7]:
-$\mathrm{BSD}(E/\mathbb{Q})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bigl{(}L(E,s),\,E(\mathbb{Q}),\,\mathrm{Reg}(E/\mathbb{Q}),\,\Omega_{E},\,c_{E/\mathbb{Q}_{p}},\,\Sha(E/\mathbb{Q})\bigr{)}.$
+Let $E/\mathbb{Q}$ be an elliptic curve defined over $\mathbb{Q}$. We define the following 6-tuple as Birch and Swinnerton-Dyer (BSD) data:
+$\mathrm{BSD}(E/\mathbb{Q})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bigl{(}L(E,s),\,E(\mathbb{Q}),\,\mathrm{Reg}(E/\mathbb{Q}),\,\Omega_{E},\,(c_{E/\mathbb{Q}_{p}})_{p},\,\Sha(E/\mathbb{Q})\bigr{)}.$
 Here, the definition of each component is as follows.
@@
   • $\Sha(E/\mathbb{Q})$: The Tate-Shafarevich group of $E/\mathbb{Q}$, conjecturally finite.
-When $E_{1}\cong E_{2}$ over $\mathbb{Q}$, we have $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$.
+When $E_{1}\cong E_{2}$ over $\mathbb{Q}$, we have $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$. Here, “=” between groups means that the groups are isomorphic as groups.
 Question.
@@
 remains invariant under isogeny under the assumption that $\Sha(E/\mathbb{Q})$ is finite ([13], Theorem 2.1), though none of individual terms in the product need be the same for isogenous curves.
-In the high-dimensional case, Jamie Bell proved the existence of pairs of abelian varieties of dimension $p-1$ (where the class number of $\mathbb{Q}(\zeta_{p})$ is not $1$, so $p\geq 23$), defined over $\mathbb{Q}$, such that over every number field they share the same BSD invariants (except for the real period), the same $n$-Selmer groups for all $n$, and the same Tate modules [2].
+In the high-dimensional case, Jamie Bell proved the existence of pairs of abelian varieties of dimension p1p-1 (where the class number of $\mathbb{Q}(\zeta_{p})$ is not $1$, so $p\geq 23$), defined over $\mathbb{Q}$, such that over every number field they share the same BSD invariants (except for the real period), the same $n$-Selmer groups for all $n$, and the same Tate modules [2].
 The pair becomes isomorphic over $\overline{\mathbb{Q}}$.
@@
 There exist infinitely many pairs $(E_{1},E_{2})$ of non-isomorphic elliptic curves over $\mathbb{Q}$ such that $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$, and the Kodaira symbols at every prime and the minimal discriminants are the same.
-As noted in Remark 5.8, it has been proved that—up to twists—there is exactly one pair of 2‑isogenous elliptic curves that share the same BSD invariants, Kodaira symbols, and minimal discriminants.
-In summary, up to quadratic twist, there exists exactly one 2‑isogenous pair of elliptic curves over $\mathbf{Q}$ that cannot be distinguished by their BSD invariants, Kodaira symbols, minimal discriminants.
+As noted in Remark 5.8, it has been proved that—up to twists—there is a unique pair of 2‑isogenous elliptic curves that share the same BSD invariants, Kodaira symbols, and minimal discriminants.
+In summary, up to quadratic twist, there exists exactly one 2‑isogenous pair of elliptic curves over $\mathbb{Q}$ that cannot be distinguished by their BSD invariants, Kodaira symbols, minimal discriminants.
 And such indistinguishable pairs occur with positive density among quadratic twists of elliptic curves satisfying the sufficient condition of Proposition 5.4.
@@
   • For an abelian group $A$ and a prime number $p$,
-    $A[p]$ :=$:=\{\,a\in A\mid pa=0\}$ ($p$‑torsion subgroup)
-    $A[p^{\infty}]$ :=$:=\{\,a\in A\mid\exists\,n\geq 1,\;p^{n}a=0\}$ ($p$‑primary component)
+    $A[p]$ $=\{\,a\in A\mid pa=0\}$ (pp‑torsion subgroup)
+    $A[p^{\infty}]$ $=\{\,a\in A\mid\exists\,n\geq 1,\;p^{n}a=0\}$ (pp‑primary component)
   • $\Delta_{E}$: minimal discriminant of elliptic curve $E$.
@@
     $E^{D}:y^{2}=x^{3}+aDx^{2}+bD^{2}x+cD^{3}.$
-  • For an isogeny $\phi$ between abelian varieties, $\hat{\phi}$ denotes the dual isogeny of $\phi$.
-  • For an elliptic curve $E/\mathbb{Q}$, another elliptic curve $E^{\prime}/\mathbb{Q}$ and a nonzero isogeny $\phi:E\to E^{\prime}$, $\mathrm{Sel}^{\phi}(E/\mathbb{Q})$ is the $\phi$-Selmer group.
-When $E=E^{\prime}$ and $\phi=[n]$ (multiplication-by-$n\geq 1$ map), we denote its Selmer group by $\mathrm{Sel}^{n}(E/\mathbb{Q})$.
+  • For an isogeny $\phi:E_{1}\to E_{2}$ between elliptic curves, $\hat{\phi}:E_{2}\to E_{1}$ denotes the dual isogeny of $\phi$.
+  • For a quadratic twist $E^{D}$ of an elliptic curve $E$ and a degree $2$ isogeny $\phi:E_{1}\to E_{2}$ defined over $\mathbb{Q}$, we also denote by $\phi$ the induced degree $2$ isogeny $\phi_{D}:E_{1}^{D}\to E_{2}^{D}$ defined over $\mathbb{Q}$.
+Accordingly, we abbreviate the $\phi_{D}$-Selmer group $\operatorname{Sel}^{\phi_{D}}(E_{1}^{D}/\mathbb{Q})$ to $\operatorname{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$.
   • $\mathrm{rank}_{an}(E/\mathbb{Q})\coloneqq\mathrm{ord}_{s=1}L(E,s)$: analytic rank of $E/\mathbb{Q}$.
@@
 The LMFDB already contains several pairs satisfying this condition, and one can easily confirm that these pairs indeed have the same BSD invariants and Kodaira symbols by Proposition 5.4.
-Applying quadratic twists then yield infinitely many further pairs; the existence of suitable twists follows in general from the Smith density theorem or classical two descent (Proposition 6.1) in the special case that we have sought.
+Applying quadratic twists then yields infinitely many further pairs; the existence of suitable twists follows in general from the Smith density theorem or classical two descent (Proposition 6.1) in the special case that we have sought.
 To align the Tate–Shafarevich groups, we simultaneously trivialize their 2-torsion subgroups. By Proposition 5.3, the entire Tate–Shafarevich groups therefore coincide.
@@
 $\lim_{H\to\infty}\frac{\#\{D\in X_{E}(D_{0},H)\,:\,\,\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{D})=r+1\}}{\#X_{E}(D_{0},H)}=\alpha_{r,u}.$
-where $\alpha_{r,u}$ denotes the limit, as $n\to\infty$, of the probability that a uniformly random $(n-u)\times n$ matrix over $\mathbb{F}_{2}$ has rank exactly $n-r$.
+where $\alpha_{r,u}$ denotes the limit, as $n\to\infty$, of the probability that a uniformly random $(n-u)\times n$ matrix over $\mathbb{F}_{2}$ has rank exactly nrn-r.
 Remark 3.3.
@@
 $\tau^{D}=\prod_{v}\left(\frac{1}{2}\,\#\,W_{v}(\phi,D)\right),$
-where $W_{v}(\phi,D)\;:=\;\ker\mskip-3.0mu\Bigl{(}H^{1}(\mathbb{Q}_{v},E_{1}^{D}[\phi])\longrightarrow H^{1}(\mathbb{Q}_{v},E_{1}^{D})\Bigr{)}$  (see Theorem 3.5 of [6]).
+where $W_{v}(\phi,D)\;:=\;\text{Ker}\ \mskip-3.0mu\Bigl{(}H^{1}(\mathbb{Q}_{v},E_{1}^{D}[\phi])\longrightarrow H^{1}(\mathbb{Q}_{v},E_{1}^{D})\Bigr{)}$  (see Theorem 3.5 of [6]).
 When $E^{D}/\mathbb{Q}$ has good reduction at an odd prime $v$, $W_{v}(\phi,D)\cong H^{1}_{\mathrm{nr}}(\mathbb{Q}_{v},E^{D}[\phi])\cong H^{1}(\widehat{\mathbb{Z}},\mathbb{Z}/2\mathbb{Z})\cong\mathbb{Z}/2\mathbb{Z}$ where $H^{1}_{\mathrm{nr}}(\mathbb{Q}_{v},E^{D}[\phi])$ is the unramified cohomology.
@@
 Summarizing the above, $\tau^{D}=\tau^{D_{0}}$.
-We use this remark to simultaneously control $\Sha(E_{1}^{D}/\mathbb{Q})$ and $\Sha(E_{2}^{D}/\mathbb{Q})$ in Section 6.
+We use this remark to simultaneously control $\Sha(E_{1}^{D}/\mathbb{Q})$ and $\Sha(E_{2}^{D}/\mathbb{Q})$ in the following Proposition.
 Proposition 3.4.
@@
 By Theorem 3.2, there exist infinitely many square-free integers $D$ such that $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$ and $e_{D}=f_{D}=1$ by taking $D_{0}=1$ and $r=0$.
-Indeed, with $D_{0}=1$, we have $u=e_{D}-f_{D}=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})=0$.
+Indeed, with $D_{0}=1$, we have $u=e_{D}-f_{D}=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}/\mathbb{Q})=0$ by Remark 3.3.
 Also, $D\in(\mathbb{Q}_{v}^{\times})^{2}$ for every $p\mid 2\Delta_{E_{1}}\Delta_{E_{2}}$ implies $\text{gcd}(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$.
@@
 For this square-free integer $D$, we have
-(3.5) #$\#\Sha(E^{D}_{i}/\mathbb{Q})[2]=\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})-\dim_{\mathbb{F}_{2}}E^{D}_{i}(\mathbb{Q})/2E^{D}_{i}(\mathbb{Q})\leq 1-1=0$
+(3.5) $\dim_{\mathbb{F}_{2}}\Sha(E^{D}_{i}/\mathbb{Q})[2]=\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})-\dim_{\mathbb{F}_{2}}E^{D}_{i}(\mathbb{Q})/2E^{D}_{i}(\mathbb{Q})\leq 1-1=0$
 for $i\in\{1,2\}$.
-Additionally, since $\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}\leq\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})\leq 2$ and #$\#\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}=2^{\mathrm{rank}(E^{D}_{i}/\mathbb{Q})+1}$, we have $\mathrm{rank}(E^{D}_{i}/\mathbb{Q})=0$ for $i\in\{1,2\}$.
+Additionally, since $\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}\leq\#\mathrm{Sel}^{2}(E^{D}_{i}/\mathbb{Q})\leq 2$ and $\dfrac{E^{D}_{i}(\mathbb{Q})}{2E^{D}_{i}(\mathbb{Q})}=2^{\mathrm{rank}(E^{D}_{i}/\mathbb{Q})+1}$, we have $\mathrm{rank}(E^{D}_{i}/\mathbb{Q})=0$ for $i\in\{1,2\}$.
 Note that the condition $\mathrm{rank}_{an}(E_{1}^{D}/\mathbb{Q})=\mathrm{rank}_{an}(E_{2}^{D}/\mathbb{Q})\equiv 0\bmod 2$ follows automatically from $\Sha(E^{D}_{1}/\mathbb{Q})[2]\cong\Sha(E^{D}_{2}/\mathbb{Q})[2]=0$ and $\mathrm{rank}(E^{D}_{1}/\mathbb{Q})=\mathrm{rank}(E^{D}_{2}/\mathbb{Q})=0$.
@@
 Theorem 4.1.
-Let $E_{1}$ and $E_{2}$ be isogenous elliptic curves over $\mathbb{Q}$, and assume $E_{1}$ has good reduction at $2$. The following are equivalent:
+Let $p$ be a prime number.
+Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous by a degree $p$ isogeny over $\mathbb{Q}$.
+Assume $E_{1}$ has good reduction at $2$.
+The following are equivalent:
   1. (1)
@@
 Proof.
-($(1)\Rightarrow(2)$) Suppose all Kodaira symbols are the same. Ogg’s formula states that
-$v_{l}(\Delta_{E})=f_{l}+m_{l}-1,$
-where $f_{l}$ is the conductor exponent and $m_{l}=\#\tilde{\epsilon}(\overline{\mathbb{F}_{l}})/\#\tilde{\epsilon}_{0}(\overline{\mathbb{F}_{l}})$.
+($(1)\Rightarrow(2)$) Ogg’s formula states that
+$v_{\ell}(\Delta_{E})=f_{\ell,E}+m_{\ell,E}-1,$
+where $f_{\ell}$ is the conductor exponent and $m_{\ell}=\#(\tilde{\epsilon}/\tilde{\epsilon}_{0})(\overline{\mathbb{F}_{\ell}})$(see [10], §11).
 Here, $\tilde{\epsilon}$ is the special fiber of the Néron minimal model of $E/\mathbb{Q}$, $\tilde{\epsilon}_{0}$ is its connected component of the identity, and $\Delta_{E}$ is the minimal discriminant of $E/\mathbb{Q}$.
-Since $E_{1}$ and $E_{2}$ are isogenous, the conductor exponent $f_{l}$ is an isogeny invariant, so $f_{l,E_{1}}=f_{l,E_{2}}$ for all primes $l$.
-Since $E_{1}$ and $E_{2}$ have the same Kodaira symbols at all primes $l$, we have $m_{l,E_{1}}=m_{l,E_{2}}$ for all primes $l$.
-Therefore, $v_{l}(\Delta_{E_{1}})=v_{l}(\Delta_{E_{2}})$ for all primes $l$.
+Since $E_{1}$ and $E_{2}$ are isogenous, the conductor exponent $f_{\ell}$ is an isogeny invariant, so $f_{\ell,E_{1}}=f_{\ell,E_{2}}$ for all primes $\ell$.
+Since $E_{1}$ and $E_{2}$ have the same Kodaira symbols at all primes $\ell$, we have $m_{\ell,E_{1}}=m_{\ell,E_{2}}$ for all primes $\ell$.
+Therefore, $v_{\ell}(\Delta_{E_{1}})=v_{\ell}(\Delta_{E_{2}})$ for all primes $\ell$.
 This implies $\Delta_{E_{1}}=\pm\Delta_{E_{2}}$.
@@
 Since $E_{1}$ and $E_{2}$ are isogenous, $f_{\ell,E_{1}}=f_{\ell,E_{2}}$, which implies $m_{\ell,E_{1}}=m_{\ell,E_{2}}$.
-By lemma 2.16 of [1], $E_{1}$ has potentially good reduction at $l>2$.
-In the additive potentially good case, the Kodaira symbol is uniquely determined by $m_{\ell}$ (See [10], IV.9, Table 4.1).
-Since both invariants are equal, the Kodaira symbols are the same.
+By Lemma 2.16 of [1], $E_{1}$ has potentially good reduction at $\ell$.
+In the additive potentially good case, the Kodaira symbol at $\ell>2$ is uniquely determined by $m_{\ell}$ (See [11], IV.9, Table 4.1).
+Since $m_{\ell,E_{1}}=m_{\ell,E_{2}}$, the Kodaira symbols are the same.
 ∎
 Theorem 4.2.
-Suppose $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ are elliptic curves such that there exists a degree $2$ isogeny between them.
-Let $l\geq 3$ be a prime and $F$ be a quadratic extension of $\mathbb{Q}_{l}$ over which $E_{1}$ (or $E_{2}$) has good reduction.
-If $\frac{\Delta_{E_{1}}}{\Delta_{E_{2}}}\in N_{F/\mathbb{Q}_{l}}$, then $c_{E_{1}/\mathbb{Q}_{l}}=c_{E_{2}/\mathbb{Q}_{l}}$.
+Suppose $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ are elliptic curves such that there exists a degree $2$ isogeny defined over $\mathbb{Q}$ between them.
+Let $\ell\geq 3$ be a prime.
+If $\Delta_{E_{1}}=\Delta_{E_{2}}$, then $c_{E_{1}/\mathbb{Q}_{\ell}}=c_{E_{2}/\mathbb{Q}_{\ell}}$.
 Proof.
-See [3], Table 1 and Theorem 6.1. ∎
-Proposition 4.3.
+Since $\Delta_{E_{1}}=\Delta_{E_{2}}$, $E_{1}$ and $E_{2}$ have potentially good reduction at $\ell$ by Lemma 2.16 of [1].
+When the Kodaira symbol at $\ell$ is not $I^{*}_{0}$, then $c_{E_{1}/\mathbb{Q}_{\ell}}=c_{E_{2}/\mathbb{Q}_{\ell}}$ (see the proof of Theorem 6.1, [3]).
+When the Kodaira symbol at $\ell$ is $I^{*}_{0}$, the result follows since $1=\frac{\Delta_{E_{1}}}{\Delta_{E_{2}}}\in N_{F/\mathbb{Q}_{\ell}}$ where $F$ is a quadratic extension of $\mathbb{Q}_{\ell}$ over which $E_{1}$ has good reduction (see [3], Table 1 and Theorem 6.1).
+Theorem 4.3.
 Let $p$ be a prime number.
-Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous by a degree $p$ isogeny over $\mathbb{Q}$.
+Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous by a degree $p$ isogeny defined over $\mathbb{Q}$.
 Assume that $E_{1}$ and $E_{2}$ have the same Tamagawa numbers at all primes and $\mathrm{rank}_{an}(E_{1}/\mathbb{Q})\equiv 0\bmod 2$.
 Then, $\Omega_{E_{1}}=\Omega_{E_{2}}$.
-In particular, if $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$ and $E_{1}$ and $E_{2}$ are isogenous by degree $p$ isogeny, then $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ have even analytic rank.
+In particular, if $\text{BSD}(E_{1}/\mathbb{Q})=\text{BSD}(E_{2}/\mathbb{Q})$ and $E_{1}$ and $E_{2}$ are isogenous by degree $p$ isogeny defined over $\mathbb{Q}$, then $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ have even analytic rank.
 Proof.
 By Theorem 8.2 of [3],
-$\Omega_{E_{1}}=\Omega_{E_{2}}\Leftrightarrow\sum_{l}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{l}}}{c_{E_{2}/\mathbb{Q}_{l}}})\equiv\text{ord}_{s=1}L(E_{1},s)\bmod 2.$
-Since Tamagawa numbers at all primes are the same, $\sum_{l}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{l}}}{c_{E_{2}/\mathbb{Q}_{l}}})=0$ holds.
+$\Omega_{E_{1}}=\Omega_{E_{2}}\Leftrightarrow\sum_{\ell}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{\ell}}}{c_{E_{2}/\mathbb{Q}_{\ell}}})\equiv\text{ord}_{s=1}L(E_{1},s)\bmod 2.$
+Since Tamagawa numbers at all primes are the same, $\sum_{l}\text{ord}_{p}(\frac{c_{E_{1}/\mathbb{Q}_{\ell}}}{c_{E_{2}/\mathbb{Q}_{\ell}}})=0$ holds.
 Thus, the equality $\Omega_{E_{1}}=\Omega_{E_{2}}$ is equivalent to the congruence $\text{ord}_{s=1}L(E_{1},s)\equiv 0\bmod 2$, which is true.
@@
 Theorem 5.1.
-If $E$ is an elliptic curve over $\mathbb{Q}$, then there are only finite number of square-free integers $D$ such that $E_{D}$ has a torsion point of order $>2$.
+Let $E$ be an elliptic curve over $\mathbb{Q}$. There are at most finitely many square-free integers $D$ for which $E^{D}(\mathbb{Q})$ contains a torsion point of order >2>2.
 Proof.
@@
 Since $\phi\circ\hat{\phi}=[p]$ and $\hat{\phi}\circ\phi=[p]$, we have $F(\phi)\circ F(\hat{\phi})=[p]$ and $F(\hat{\phi})\circ F(\phi)=[p]$.
-For a prime $l\neq p$ and for all $m$, since $\gcd(p,l)=1$, $F([p])=p:F(E_{1})[l^{\infty}]\to F(E_{2})[l^{\infty}]$ is bijective, it follows that $F(\phi):F(E_{1})[l^{\infty}]\to F(E_{2})[l^{\infty}]$ is an isomorphism.
+For a prime $\ell\neq p$ and for all $m$, since $\gcd(p,\ell)=1$, $F([p])=p:F(E_{1})[\ell^{\infty}]\to F(E_{1})[\ell^{\infty}]$ is bijective, it follows that $F(\phi):F(E_{1})[\ell^{\infty}]\to F(E_{2})[\ell^{\infty}]$ is an isomorphism.
 Since $F(E_{1})$ and $F(E_{2})$ are torsion groups, we have
@@
 Proposition 5.4.
-Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous via a degree–$2$ isogeny defined over $\mathbb{Q}$. Assume that
+Let $E_{1}/\mathbb{Q}$ and $E_{2}/\mathbb{Q}$ be elliptic curves that are isogenous via a degree $2$ isogeny defined over $\mathbb{Q}$. Assume that
   1. (1)
@@
 By Theorem 4.1, the Kodaira symbols are the same.
-By Theorem 4.2, Tamagawa numbers at $l\geq 3$ are the same because $\Delta_{E_{1}}=\Delta_{E_{2}}$.
-For the period, $\Omega_{E_{1}}=\Omega_{E_{2}}$ follows from Proposition 4.3.
+By Theorem 4.2, Tamagawa numbers at $\ell\geq 3$ are the same because $\Delta_{E_{1}}=\Delta_{E_{2}}$.
+For the period, $\Omega_{E_{1}}=\Omega_{E_{2}}$ follows from Theorem 4.3.
 Because $\mathrm{rank}(E_{i}/\mathbb{Q})=0$, $\mathrm{Reg}(E_{i}/\mathbb{Q})=1$ for $i\in\{1,2\}$.
@@
 Example 5.5.
-Examples of pairs $(E_{a},E_{b})$ of elliptic curves that satisfy conditions (1), (2), and (3) of Proposition 5.4 are $(38025.ck1,38025.ck2)$ (this is the same as the pair described in Proposition 6.1), $(38025.t1,38025.t2)$, $(38025.u1,38025.u2)$, $(38025.cl1,38025.cl2)$, $(207025.bf1,207025.bf2)$, $(207025.bu1,207025.bu2)$, and $(207025.cj1,207025.cj2)$.
-As explained in Remark 5.8, these are unique pairs of 2-isogenous elliptic curves that share the same minimal discriminants, up to twists.
+An example of the pair $(E_{1},E_{2})$ that satisfies the conditions (1), (2), and (3) is $(y^{2}=x^{3}+25350x^{2}+2471625x,y^{2}=x^{3}-50700x^{2}+632736000x)$.
+The LMFDB labels of $E_{1}$ and $E_{2}$ are 38025.ck1 and 38025.ck2, respectively.
+Indeed, for $(1)$, $E_{1}(\mathbb{Q})[2^{\infty}]\cong E_{2}(\mathbb{Q})[2^{\infty}]$.
+For $(2)$, $\Sha(E_{1}/\mathbb{Q})[2]=\Sha(E_{2}/\mathbb{Q})[2]=0$ and thus $\Sha(E_{1}/\mathbb{Q})[2^{\infty}]\cong\Sha(E_{2}/\mathbb{Q})[2^{\infty}]=0$.
+For $(3)$, $\Delta_{E_{1}}=\Delta_{E_{2}}=3^{6}\cdot 5^{9}\cdot 13^{9}$.
+As explained in Remark 5.8, this is a unique pair of 2 isogenous elliptic curves that shares the same minimal discriminants, up to twists.
 Theorem 5.6.
-Suppose that $E_{1}(\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}$ and $E_{1}/\mathbb{Q}$ has good reduction at $2$.
-Suppose $(E_{1},E_{2})$ is a pair of a balanced isogenous elliptic curves with $E_{1}(\mathbb{Q})[2^{\infty}]\cong E_{2}(\mathbb{Q})[2^{\infty}]$, and $\Delta_{E_{1}}=\Delta_{E_{2}}$.
-Additionally, suppose $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}/\mathbb{Q})$.
-Then there exist infinitely many square-free integers $D$ such that $\text{BSD}(E^{D}_{1}/\mathbb{Q})=\text{BSD}(E^{D}_{2}/\mathbb{Q})$ and Kodaira symbols at every prime and minimal discriminants are the same.
+Suppose $E_{i}/\mathbb{Q}\ (i\in\{1,2\})$ is an elliptic curve with $E_{i}(\mathbb{Q})[2]\simeq\mathbb{Z}/2\mathbb{Z}$.
+Suppose that $E_{1}$ and $E_{2}$ are isogenous via a degree $2$ balanced isogeny $\phi$.
+Suppose that $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}/\mathbb{Q})=\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}/\mathbb{Q})$.
+Also, suppose that $E_{1}$ has good reduction at $2$ and $\Delta_{E_{1}}=\Delta_{E_{2}}$.
+Then, there exist infinitely many square-free integers $D$ such that $\text{BSD}(E^{D}_{1}/\mathbb{Q})=\text{BSD}(E^{D}_{2}/\mathbb{Q})$ and Kodaira symbols at every prime and minimal discriminants are the same.
 Proof.
@@
 For condition (1), For a square-free integer $D$, $E_{1}^{D}(\mathbb{Q})[2]\cong E_{2}^{D}(\mathbb{Q})[2]$.
-By Theorem 5.1, there are finitely many $D$ such that $E^{D}_{1}(\mathbb{Q})_{\text{tor}}$ or $E^{D}_{2}(\mathbb{Q})_{\text{tor}}$ acquire points of order greater than 4.
+By Theorem 5.1, there are finitely many $D$ such that $E^{D}_{1}(\mathbb{Q})_{\text{tor}}$ or $E^{D}_{2}(\mathbb{Q})_{\text{tor}}$ acquire points of order greater than 2.
 Thus, for almost all square-free integers $D$, we have $E^{D}_{1}(\mathbb{Q})[2^{\infty}]\cong E^{D}_{2}(\mathbb{Q})[2^{\infty}]$.
@@
 For (3), let us take $(D,2\Delta_{E_{1}}\Delta_{E_{2}})=1$.
-Note that $E^{D}_{1}/\mathbb{Q},E^{D}_{2}/\mathbb{Q}$ is global minimal when the model of $E_{1}/\mathbb{Q},E_{2}/\mathbb{Q}$ is global minimal because $(D,\Delta_{E_{1}}\Delta_{E_{2}})=1$.
+Note that $E^{D}_{1}/\mathbb{Q},E^{D}_{2}/\mathbb{Q}$ are global minimal when the model of $E_{1}/\mathbb{Q},E_{2}/\mathbb{Q}$ are global minimal because $(D,\Delta_{E_{1}}\Delta_{E_{2}})=1$.
 Also, $E^{D}_{1}$ has good reduction at $2$.
@@
 Proof.
-The pairs in Example 5.5 satisfy the hypothesis of Proposition 5.4. By Theorem 5.6, infinitely many pairs $(E^{D}_{a},E^{D}_{b})$ are the pairs we have sought. ∎
+The pairs in Example 5.5 satisfy the hypothesis of Proposition 5.4 and Theorem 5.6. By Theorem 5.6, infinitely many pairs $(E^{D}_{1},E^{D}_{2})$ are the pairs we have sought. ∎
 Finally, we see that these examples are very rare.
@@
 Recent work [1] proved that there is only one pair of discriminant twins among 2-isogenous curves up to twists: $(4225.h1,4225.h2)$.
-The curve 4225.h1 is a twist of our 38025.ck1 by $-195$.
+The curve 4225.h1 is a twist of our 38025.ck1 by 195-195.
 For distinct square–free integer $D_{1},D_{2}$, $E^{D_{1}},E^{D_{2}}$ do not have the same BSD invariants since the real period $\Omega_{E^{D_{1}}}=\frac{\Omega_{E}}{\sqrt{|D_{1}|}},\Omega_{E^{D_{2}}}=\frac{\Omega_{E}}{\sqrt{|D_{2}|}}$ never coincide.
@@
 In this chapter, independently of Smith’s results, we give a concrete infinite family using classical 2-descent (see for example, [11], X, Remark 3.7 and Proposition 4.9) with a balanced isogeny case.
-The key point is that since this is a balanced isogeny case, we can simultaneously control $\Sha(E_{1}/\mathbb{Q})$ and $\Sha(E_{2}/\mathbb{Q})$ by using Remark 3.3.
 6.1. Two descent for balanced isogenous pairs of elliptic curves
@@
 Let $E_{1}:y^{2}=x^{3}+25350x^{2}+2471625x$ and $E_{2}:y^{2}=x^{3}-50700x^{2}+632736000x$ (labeled as (38025.ck1, 38025.ck2) in LMFDB).
-Take a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$.
+Take a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$.
 For such $D$, $\Sha(E_{1}^{D}/\mathbb{Q})[2]=0$ and $\Sha(E_{2}^{D}/\mathbb{Q})[2]=0$.
@@
 Let us define
-$S\coloneqq\{v\mid v\text{ divides }\deg\phi\}\cup\{v\mid v\text{ divides }\Delta_{E_{1}}\Delta_{E_{2}}\}\cup\{\infty\}=\{2,3,5,13,D,\infty\}$
+$S\coloneqq\{v\mid v\text{ divides }\deg\phi\}\cup\{v\mid v\text{ divides }\Delta_{E^{D}_{1}}\Delta_{E^{D}_{2}}\}\cup\{\infty\}=\{2,3,5,13,D,\infty\}$
 and $\mathbb{Q}(S,2)\coloneqq\{\bar{d}\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}\mid v(d)\equiv 0\pmod{2}\text{ for all }v\notin S\}$. Then,
-#$\#\mathbb{Q}(S,2)=\#\{(-1)^{n_{0}}2^{n_{1}}3^{n_{2}}5^{n_{3}}13^{n_{4}}D^{n_{5}}\mid 0\leq n_{1},n_{2},n_{3},n_{4},n_{5}\leq 1\}=64.$
+$\mathbb{Q}(S,2)=\#\{(-1)^{n_{0}}2^{n_{1}}3^{n_{2}}5^{n_{3}}13^{n_{4}}D^{n_{5}}\mid 0\leq n_{1},n_{2},n_{3},n_{4},n_{5}\leq 1\}=64.$
 Via the following isomorphism, the Selmer group can be calculated by checking the local solubility of torsors $C_{d}$.
-$\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})\cong\left\{\bar{d}\in\mathbb{Q}(S,2)\;\middle|\;\begin{aligned} &\bar{C}_{d}(\mathbb{Q}_{v})\neq\emptyset,\forall v\in S,\text{where }\ \bar{C}_{d}\ \text{is the projective closure of}\\ &\text{affine curve}\ C_{d}:dy^{2}=d^{2}-50700dDx^{2}+632736000D^{2}x^{4}\ \text{in}\ \mathbb{P}^{3}\end{aligned}\right\}$
+$\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})\cong\left\{\bar{d}\in\mathbb{Q}(S,2)\;\middle|\;\begin{aligned} &\bar{C}_{d}(\mathbb{Q}_{v})\neq\emptyset,\forall v\in S,\text{where }\ \bar{C}_{d}\ \text{is the projective closure of}\\ &\text{affine curve}\ C_{d}:dy^{2}=d^{2}-50700dDx^{2}+632736000D^{2}x^{4}\ \text{in}\ \mathbb{P}^{3}\end{aligned}\right\}$
 (see [11], X, Remark 3.7 and Proposition 4.9). Here, $50700=2^{2}\times 3\times 5^{2}\times 13^{2},632736000=2^{8}\times 3^{2}\times 5^{3}\times 13^{3}$.
 $C_{d}$ has two points at infinity $\infty\coloneqq[0,0,\pm\sqrt{\frac{632736000}{d}},1]$ in $\mathbb{P}^{3}$ (see [11], X, Remark 3.7). Since $\sqrt{\frac{632736000}{65}}=3120$, we note that
-(6.2) $65\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q}).$
+(6.2) $65\in\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q}).$
 For $64$ candidates $d\in\mathbb{Q}(S,2)$, we check which $d$ is in $\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$.
@@
 $\bar{C}_{d}$ $p(d)$ third coordinate of points at $\infty$
-$\bar{C}_{d}\,(d<0)$infty ±$\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{R}\ \text{since}\ d<0$
-$\bar{C}_{d}\,(3\mid d)$ 3 ±$\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{Q}_{3}\ \text{since}\ v_{3}(\frac{632736000}{d})=2-n_{2}=1$
-$\bar{C}_{2}$ 5 ±$\pm\sqrt{\frac{632736000}{2}}\notin\mathbb{Q}_{5}$
-$\bar{C}_{13}$ 5 ±$\pm\sqrt{\frac{632736000}{13}}\notin\mathbb{Q}_{5}$
-$\bar{C}_{D}$ 5 ±$\pm\sqrt{\frac{632736000}{D}}\notin\mathbb{Q}_{5}$
-$\bar{C}_{5D}$ 13 ±$\pm\sqrt{\frac{632736000}{5D}}\notin\mathbb{Q}_{13}$
-$\bar{C}_{2D}$ 13 ±$\pm\sqrt{\frac{632736000}{2D}}\notin\mathbb{Q}_{13}$
+$\bar{C}_{d}\,(d<0)$ $\infty$ $\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{R}\ \text{since}\ d<0$
+$\bar{C}_{d}\,(3\mid d)$ 3 $\pm\sqrt{\frac{632736000}{d}}\notin\mathbb{Q}_{3}\ \text{since}\ v_{3}(\frac{632736000}{d})=2-n_{2}=1$
+$\bar{C}_{2}$ 5 $\pm\sqrt{\frac{632736000}{2}}\notin\mathbb{Q}_{5}$
+$\bar{C}_{13}$ 5 $\pm\sqrt{\frac{632736000}{13}}\notin\mathbb{Q}_{5}$
+$\bar{C}_{D}$ 5 $\pm\sqrt{\frac{632736000}{D}}\notin\mathbb{Q}_{5}$
+$\bar{C}_{5D}$ 13 $\pm\sqrt{\frac{632736000}{5D}}\notin\mathbb{Q}_{13}$
+$\bar{C}_{2D}$ 13 $\pm\sqrt{\frac{632736000}{2D}}\notin\mathbb{Q}_{13}$
 $\bar{C}_{10D}$ $D$ $\pm\sqrt{\frac{63273600}{D}}\notin\mathbb{Q}_{D}$
-$\bar{C}_{10}$ 2 ±$\pm\sqrt{\frac{632736000}{10}}\notin\mathbb{Q}_{2}$
-When $d<0$, $\bar{C}_{d}\notin\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$.
+$\bar{C}_{10}$ 2 $\pm\sqrt{\frac{632736000}{10}}\notin\mathbb{Q}_{2}$
+When $d<0$, $\bar{C}_{d}\notin\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$.
 Thus, $n_{0}=0$.
@@
 Now, $\frac{f(x)}{9}\in\mathbb{Z}_{3}$ and $\frac{f(x)}{9}=d^{\prime 2}-16900Dd^{\prime}x^{2}+70304000D^{2}x^{4}\equiv d^{\prime 2}+2Dd^{\prime}x^{2}+2D^{2}x^{2}\equiv 1+2Dd^{\prime}+2D^{2}\not\equiv 0\bmod 3$, so $v_{3}(f(x))=2$.
-Thus, $n_{2}$ should be 00.
-So, the candidate of $d\in\mathbb{Q}(S,2)$ such that $d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$ is of the form $2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$.
+Thus, $n_{2}$ should be 0.
+So, the candidate of $d\in\mathbb{Q}(S,2)$ such that $d\in\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$ is of the form $2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$.
 Thus, $d=2^{n_{1}}5^{n_{3}}13^{n_{4}}D^{n_{5}}$.
-The left candidates of $d$ is $d=2,5,13,65,130,D,5D,13D,65D,10,26,10D,26D,130D$.
-To check these candidates, it is sufficient to prove $d=2,13,D,5D,2D,10D,10$ does not belong to $\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$. Indeed,
-$d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})\Leftrightarrow 65d\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$
-holds since $65\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$(see (6.2)).
-By the following Lemma 6.3, we have $\text{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})=\{1,65\}\cong\mathbb{Z}/2\mathbb{Z}.$
-2-descent shows that $\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\phi}(E_{1}^{17}/\mathbb{Q})-\mathrm{dim}_{\mathbb{F}_{2}}\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})=0$.
-In Remark 6.4, let us take $D_{0}=17$ so that $17D\in{\mathbb{Q}_{v}^{\times}}^{2}$.
-This condition is equivalent to $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, and $D>0$.
-Then, by Remark 6.4, since $\phi$ is balanced, we have $\mathrm{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}$.
+The left candidates of $d$ is d=2,5,13,65,130,D,5D,13D,65D,10,26,10D,26D,130Dd=2,5,13,65,130,D,5D,13D,65D,10,26,10D,26D,130D.
+To check these candidates, it is sufficient to prove $d=2,13,D,5D,2D,10D,10$ does not belong to $\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$. Indeed,
+$d\in\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})\Leftrightarrow 65d\in\mathrm{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})$
+holds since $65\in\mathrm{Sel}^{\phi}(E^{D}/\mathbb{Q})$(see (6.2)). By the following Lemma 6.3, we have $\text{Sel}^{\phi}(E_{1}^{D}/\mathbb{Q})=\{1,65\}\cong\mathbb{Z}/2\mathbb{Z}.$
+By Remark 3.3, $\tau^{D}=\prod_{p}\frac{1}{2}\#W_{p}(\phi,D)$.
+For any prime $p\neq 2$, by Lemma 2.1 (ii) of [14], we have $\frac{1}{2}\#W_{p}(\phi,D)=\frac{c_{E_{1}/\mathbb{Q}_{p}}}{c_{E_{2}/\mathbb{Q}_{p}}}=1$ where the last equality holds by Theorem 4.2.
+For $p=2$, $\frac{1}{2}{\#W_{2}(\phi,D)}=1$ since $E^{D}/\mathbb{Q}$ has good reduction at $2$.
+At the infinite place $p=\infty$, by Lemma 2.1 (iii), we have $\frac{1}{2}\#W_{\infty}(\phi,D)=1$ since $2471625D^{2}>0$.
+Thus, we have $\text{dim}_{\mathbb{F}_{2}}\text{Sel}^{\phi}(E^{D}_{1}/\mathbb{Q})=\text{dim}_{\mathbb{F}_{2}}\text{Sel}^{\hat{\phi}}(E_{2}^{D}/\mathbb{Q})$ since $\tau^{D}=1$.
 By (3.5) in the proof of Proposition 3.4, we have $\Sha(E^{D}_{1}/\mathbb{Q})[2]=\Sha(E^{D}_{2}/\mathbb{Q})[2]=0$ and $\mathrm{rank}(E_{1}^{D}/\mathbb{Q})=\mathrm{rank}(E_{2}^{D}/\mathbb{Q})=0$ for infinitely many square-free integer $D$.
@@
 When $v_{5}(x)<0$, $v_{5}(f_{D}(x))=\text{min}\{0,2+2v_{5}(x),3+4v_{5}(x)\}=3+4v_{5}(x)$ is odd and thus $C_{D}(\mathbb{Q}_{5})=\emptyset$.
-[…]
+For $d=5D$, take $p(5D)=13$.
+$C_{5D}:y^{2}=5D-50\,700D\,x^{2}+126\,547\,200D\,x^{4}$.
+When $v_{13}(x)\geq 0$, $y^{2}\equiv 5D\bmod 13$.
+This contradicts that $\left(\frac{5D}{13}\right)=-\left(\frac{D}{13}\right)=-1$.
+When $v_{13}(x)<0$, since $v_{13}(f_{5D}(x))=\text{min}\{0,2+2v_{13}(x),3+4v_{13}(x)\}=3+4v_{13}(x)$ is odd, we have $C_{5D}(\mathbb{Q}_{13})=\emptyset$.
 For $d=2D$, take $p(2D)=13$.
@@
 When $v_{D}(x)=0$, we claim $v_{D}(g(x))=0$ where $g(x)\coloneqq 10-50700x^{2}+63273600x^{4}$.
-The discriminant $\Delta$ of the polynomial $1063273600t^{2}-50700t+$ 10 is $\Delta=2^{4}\times 3^{2}\times 5^{3}\times 13^{3}$.
+The discriminant $\Delta$ of the polynomial $63273600t^{2}-50700t+10$ is $\Delta=2^{4}\times 3^{2}\times 5^{3}\times 13^{3}$.
 Since $\left(\frac{\Delta}{D}\right)=\left(\frac{5\times 13}{D}\right)=\left(\frac{5}{D}\right)\left(\frac{13}{D}\right)=-1$, $v_{D}(g(x))=0$ because $g(x)\equiv 0\bmod D$ does not have solutions.
@@
 When $v_{2}(x)=-2$, by substituting $W=4x$, $C_{10}:4y^{2}=40-12675DW^{2}+988650D^{2}W^{4}$.
-Since $4y^{2}\equiv 0-3\times 1\times 1+2\equiv 7\bmod 8$, $v_{2}(x)$ is not $-2$.
+Since $4y^{2}\equiv 0-3\times 1\times 1+2\equiv 7\bmod 8$, $v_{2}(x)$ is not 2-2.
 The left case is $v_{2}(x)\leq-2$ or $v_{2}(x)\geq 0$.
@@
 6.2. Pairs of elliptic curves sharing the BSD invariants, Kodaira symbols and minimal discriminants
-Combining Proposition 5.4 and Proposition 6.1, for a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{3}\right)=-1$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$, and $D$ is large enough so that $E_{1}^{D}(\mathbb{Q})_{\text{tor}}$ and $E_{2}^{D}(\mathbb{Q})_{\text{tor}}$ do not have any elements of order $>2$, we have $\text{BSD}(E_{1}^{D}/\mathbb{Q})=\text{BSD}(E_{2}^{D}/\mathbb{Q})$ and the Kodaira symbols of $E_{1}$ and $E_{2}$ are the same.
+Combining Proposition 5.4 and Proposition 6.1, for a prime number $D>0$ such that $D\equiv 1\pmod{8}$, $\left(\frac{D}{5}\right)=-1$, $\left(\frac{D}{13}\right)=1$, $\gcd(D,2\cdot 3\cdot 5\cdot 13)=1$, and $D$ is large enough so that $E_{1}^{D}(\mathbb{Q})_{\text{tor}}$ and $E_{2}^{D}(\mathbb{Q})_{\text{tor}}$ do not have any elements of order >2>2, we have $\text{BSD}(E_{1}^{D}/\mathbb{Q})=\text{BSD}(E_{2}^{D}/\mathbb{Q})$ and the Kodaira symbols of $E_{1}$ and $E_{2}$ are the same.
 We provide a table of pairs of non-isomorphic elliptic curves that share the same BSD invariants.
@@
 Regulator 1 1
-Real period $\dfrac{0.209\cdots}{\sqrt{|D|}}$ $\dfrac{0.209\cdots}{\sqrt{|D|}}$
+Real period $\dfrac{0.209\cdots}{\sqrt{D}}$ $\dfrac{0.209\cdots}{\sqrt{D}}$
 Tamagawa number $2(p=3),2(p=5),2(p=13)$ $2(p=3),2(p=5),2(p=13)$,
@@
 $\Sha(E/\mathbb{Q})[2^{\infty}]$ 0 0
-$\Sha(E/\mathbb{Q})$dagger $\dagger$
-For $\dagger$, note that $\Sha(E/\mathbb{Q})[l^{\infty}]$ remains unknown, but they agree as abelian groups.
+$\Sha(E/\mathbb{Q})$ $\dagger$ $\dagger$
+For $\dagger$, note that $\Sha(E/\mathbb{Q})[\ell^{\infty}]$ remains unknown, but they agree as abelian groups by Proposition 5.3 (2).
 Theshaded areas in the table indicate the conditions from Proposition 5.4 and this ensures all BSD invariants and the Kodaira symbol coincide.